OpenStudy (christos):

How to find "m" and "c" in an equation of a line if you know x and y? y = mx + c

OpenStudy (anonymous):

when you have 2 unknowns, you will need two types of information. case 1) when two points are given, first find "m" then plug it in the above equation. then use any of those two points to get the value of "b"

OpenStudy (anonymous):

case 2) information concerning another line. like line is parallel or normal gives you "m" directly then use it like above case.

OpenStudy (anonymous):

if you have a particular question, go ahead

OpenStudy (christos):

But c and m are not known to me so I have 2 different uknown numbers how do I find them with only y and x to be known

OpenStudy (anonymous):

ok. lets take an example. ask a particular problem.

OpenStudy (christos):

4 = m3 + c what's m and c?

hartnn (hartnn):

i would say, you simply cannot find m and c you'll need 2 pairs of points (2 values of x and y each) to get the value of m and c.

OpenStudy (christos):

(4-2)=m(6-3) + c what is it now?

OpenStudy (christos):

Was that the tow pairs you were talking about?

hartnn (hartnn):

so, 2 points you considered are 4,6 and 2,3 ??

OpenStudy (christos):

yeah

hartnn (hartnn):

then you have 2 equations and 2 unknowns, you can find m and c 1st equation, using the point 4,6 6 = 4m+c can you form 2nd equation ?

OpenStudy (christos):

3 = -2m + c ?

hartnn (hartnn):

why -2 ? x=+2, right ? so in y=mx+c 3=2m+c right ?

OpenStudy (christos):

Yes true that i am sorry

hartnn (hartnn):

so, 2 equations, 2 unkowns, will give you m and c, can you find them now ?

OpenStudy (christos):

3=2m+c 6 = 4m+c 6-4m=c 3 = 2m+6-4m

hartnn (hartnn):

good, go ahead.

hartnn (hartnn):

you'll be able to find 'm' easily from that. then just plug in value of m in any of the equation to get c

OpenStudy (christos):

So if I find the value of c and m like this it will be the same value of m and c for the whole aequation meaning this (4-2)=m(6-3) + c ?

hartnn (hartnn):

firstly, you've interchanges x and y shouldn't it be (6-3) = (4-2)m+c ?? and yes the values of m and c would be the same because you have subtracted the equation. nevertheless you won't be able to find m and c using just that 1 equation...